On-the-Go, Inc., produces two models of traveling cases for laptop computers: the Programmer and the Executive. The bags have the following characteristics:

Programmer Executive
Selling price per bag $ 60 $ 90
Variable cost per bag $ 30 $ 30
Expected sales (bags) per year 8,000 12,000

The total fixed costs per year for the company are $665,000.

Required:

a. What is the anticipated level of profits for the expected sales volumes?

b. Assuming that the product mix is the same at the break-even point, compute the break-even point. (Round your final answer up to the nearest whole unit.)

c. If the product sales mix were to change to nine Programmer-style bags for each Executive-style bag, what would be the new break-even volume for On-the-Go? (Round your final answer up to the nearest whole unit.)

References
Answer #1

Answers

  • All working forms part of the answer
  • Requirement ‘a’

Programmer

Executive

Total

Sale price per bag

$60

$90

Variable cost per bag

$30

$30

Contribution margin per bag

$30

$60

No. of bags

                      8,000

                  12,000

Total Contribution margin

$240,000

$720,000

$960,000

Total Fixed Cost

$665,000

Net Income = Profits

$295,000

  • Requirement ‘b’

--Working

Working

Programmer

Executive

A

Sale price per pag

$60

$90

B

Variable cost per bag

$30

$30

C = A - B

Contribution margin per bag

$30

$60

D

No. of bags

                      8,000

                  12,000

E = D/20000 bags

Sales Mix %

40%

60%

F = C x E

Weighted Average contribution margin

$                   12.00

$                 36.00

--Answer

A

Total Fixed Cost

$        665,000.00

B = $12 + $ 36

Weighted Average contribution margin

$                   48.00

C = A/B

Break Even point in no. of bags

                    13,854

  • Requirement ‘c’

--Working

Working

Programmer

Executive

A

Sale price per pag

$60

$90

B

Variable cost per bag

$30

$30

C = A - B

Contribution margin per bag

$30

$60

D

No. of bags

                              9

                             1

E = D/10

Sales Mix %

90%

1%

F = C x E

Weighted Average contribution margin

$                   27.00

$                    0.60

--Answer

A

Total Fixed Cost

$        665,000.00

B = 27 + 0.6

Weighted Average contribution margin

$                   27.60

C = A/B

Break Even point in no. of bags

                    24,094